OpenStudy (anonymous):

Integrate cost/1+sint dt

OpenStudy (kainui):

Don't bother with the cross multiplication. Just do a substitution. u=1+sint du=cost

OpenStudy (ikram002p):

huh lolz yeah xD

OpenStudy (anonymous):

I need to mow the lawn >.<

OpenStudy (ikram002p):

lolz :3

OpenStudy (anonymous):

\[ \int \frac{\cos t}{1+\sin t}dt = \int \frac{d(1+\sin t)}{1+\sin t} \]